### v9

# Densities Of Primes And Primitive Roots

Abstract: Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers.
Mathematics Subject Classifications: Primary 11A07, Secondary 11N37.
Keywords: Prime Number; Primitive Root; Artin Primitive Root Conjecture.
1 Introduction
The symbol $\mathbb{Z}=\{0,\pm 1,\pm 2,\ldots\}$ denotes the set of integers, and the symbol $\mathbb{P}=\{2,3,5,\ldots\}$ denotes the set of prime numbers. The constant $\delta(u)\geq 0$ is the density of the subset of primes
$\mathcal{P}_{u}=\left\{p\in\mathbb{P}:\operatorname{ord}_{p}(u)=p-1\right\}\subset\mathbb{P}$ (1)
with a fixed primitive root $u\in\mathbb{Z}$. Let $u\neq\pm 1,v^{2}$ be a fixed integer, and let $x\geq 1$ be a large number. The expected number of primes $p\leq x$ with a fixed primitive root $u\operatorname{mod}p$ has the asymptotic formula
$\pi_{u}(x)=\#\left\{p\leqslant x:\operatorname{ord}_{p}(u)=p-1\right\}=\delta(u)\operatorname{li}(x)+O\left(x(\log x)^{-2}\right),$ (2)
where $\operatorname{li}(x)$ is the logarithm integral, as $x\rightarrow\infty$.
A conditional proof of this result was achieved in [23], and simplified sketches of the proof appear in [30, p. 8], and similar references. The determination of the constant $\delta(u)\geq 0$ for a fixed integer $u\in\mathbb{Z}$ is an interesting technical problem, [23, p. 218], [27], [28], et alii. An introduction to its historical development, and its calculations is covered in [30, pp. 3–10], and [42].
The Artin primitive root conjecture on average
$x^{-1}\sum_{u\leq x}\pi_{u}(x)=a_{0}\operatorname{li}(x)+O\left(x(\log x)^{-B}\right),$ (3)
where $a_{0}=\prod_{p\geq 2}\left(1-p^{-1}(p-1)^{-1}\right)$ is Artin constant, and $B>1$ is an arbitrary number, was proved in [18] unconditionally, and refined in [42]. These works had shown that almost all admissible integers $u\in\mathbb{Z}-\left\{-1,1,v^{2}:v\in\mathbb{Z}\right\}$ are primitive roots for infinitely many primes. The number of exceptions is a subset of zero density in $\mathbb{Z}$. The individual quantity $\pi_{u}(x)$ in (2) can be slightly different from the average quantity in (3). The variations, discovered by the Lehmers using numerical experiments, depend on the primes decomposition of the fixed value $u$, see [23, p. 220], [30, p. 3], and similar references for the exact formula for the density $\delta(u)\geq 0$.
The Artin primitive root conjecture for functions fields was proved by Bilharz, see [40, Theorem 10.11], and the same conjecture for polynomials over finite fields was proved in [38]. However, there is no known infinite sequence of primes with a fixed primitive root. The current literature has results infinite sequences of primes with unknown primitive root in a small finite set. For example, in [19] it was proved that for a fixed primes triple $q>r>s\geq 2$, the subset of integers
$\mathcal{A}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 3\right\},$ (4)
contains a primitive root for infinitely many primes. This result was later reduced to the smaller subset
$\mathcal{B}(q,r,s)=\left\{q^{a}r^{b}s^{c}:0\leq a,b,c\leq 1\right\},$ (5)
see [22]. In both of these results, the lower bound for the number of primes $p\leq x$ with a fixed primitive root in either of the subset $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$ has the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$ for all large number $x\geq 1$.
These results have been extended to quadratic numbers fields in [8], [36], [41], and most recently in [1]. Other related results are given in [31], [16], [37], et alii.
The technique explored in this note provides an improved lower bound for the number of primes $p\leq x$ with a fixed primitive root $u\operatorname{mod}p$ for all large number $x\geq 1$. And the fixed primitive root $u$ is not restricted to a small finite subset such as $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$.
Theorem 1.1.
A fixed integer $u\neq\pm 1,v^{2}$ is a primitive root mod $p$ for infinitely many primes $p\geq 2$. In addition, the density of these primes satisfies
$\pi_{u}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(u)=p-1\right\}=\delta(u)\operatorname{li}(x)+O\left(x(\log x)^{-2}\right),$ (6)
where $\operatorname{li}(x)$ is the logarithm integral, and $\delta(u)>0$ is a constant, for all large numbers $x\geq 1$.
Sections 2 to 6 introduce the notation and some standard results related to or applicable to the investigation of primitive roots in cyclic groups. The last section presents a proof of Theorem 1.1.
2 Exponentials and Character Sums
A few standard definitions and other basic results in the theory of exponential and character sums are reviewed in this section. All the estimates are unconditional.
2.1 Simple Characters Sums
Let $G$ be a finite group of order q=#Gq=\#G. The order $\operatorname{ord}(u)$ of an element $u\in G$ is the smallest integer $d\mid q$ such that $u^{d}=1$. An element $\tau\in G$ is called a primitive element if it has order $\operatorname{ord}(\tau)=q$. A cyclic group $G$ is a group generated by a primitive element $\tau\in G$. Given a primitive root $\tau\in G$, every element $0\neq u\in G$ in a cyclic group has a representation as $u=\tau^{v},0\leq v<q$. The integer $v=\log u$ is called the discrete logarithm of $u\neq 0$ with respect to $\tau$.
A character $\chi$ modulo $q\geq 2$, is a complex-valued periodic function $\chi:\mathbb{N}\longrightarrow\mathbb{C}$, and it has order $\operatorname{ord}(\chi)=d\geqslant 1$ if and only if $\chi$$(n)^{d}=1$ for all integers $n\in\mathbb{N},\gcd(n,q)=1$. For $q\neq 2^{r},r\geq 2$, a multiplicative character $\chi$ of order $\operatorname{ord}(\chi)=d\mid q$ has a representation as
$\chi(u)=e^{i2\pi k\log u/(p-1)},$ (7)
where $v=\log u$ is the discrete logarithm of $u\neq 0$ with respect to some primitive root, and for some $k\geq 1$, see [29, p. 187], [34, p. 118], and [25, p. 271] for more details.
Lemma 2.1.
For a fixed integer $u\neq 0$, and an integer $q\in\mathbb{N}$, let $\chi\neq 1$ be nonprincipal character mod $q$, then
[…]
An additive character $\psi$ of order $\operatorname{ord}(\psi)=q$ has a representation as
$\psi(n)=e^{i2\pi kn/q},$ (8)
for some $k\geq 1,\gcd(k,q)=1$, see [29, p. 187], [34, p. 118], and [25, p. 271]. The additive character sums are quite similar to Lemma 2.1.
Lemma 2.2.
For a fixed integer $u$, and an integer $q\in\mathbb{N}$, let $\psi$ be an additive character of order $\operatorname{ord}\psi=q$, then
[…]
3 Representations of the Characteristic Functions
he characteristic function $\Psi:G\longrightarrow\{0,1\}$ of primitive elements is one of the standard analytic tools employed to investigate the various properties of primitive roots in cyclic groups $G$. Many equivalent representations of the characteristic function $\Psi$ of primitive elements are possible. Several of these representations are studied in this section.
3.1 Primitive Roots Tests
For a prime $p\geq 2$, the multiplicative group of the finite fields $\mathbb{F}_{p}$ is a cyclic group for all primes.
Definition 3.1.
The order $\min\{k\in\mathbb{N}:u^{k}\equiv 1\bmod p\}$ of an element $u\in\mathbb{F}_{p}$ is denoted by $\operatorname{ord}_{p}(u)$. An element is a primitive root if and only if $\operatorname{ord}_{p}(u)=p-1$.
Lemma 3.1.
Let $u\in\mathbb{F}_{p}$. Then, $u$ is a primitive ro0t if and only if $u^{(p-1)/q}\not\equiv 1\bmod p$ for all prime divisors $q\mid p-1$.
Basically, this is the classical Lucas-Lehmer primitive root test on the cyclic group $\mathbb{F}_{p}$.
3.2 Divisors Dependent Characteristic Function
A representation of the characteristic function dependent on the orders of the cyclic groups is given below. This representation is sensitive to the primes decompositions $q=p_{1}^{e_{1}}p_{2}^{e_{2}}\cdots p_{t}^{e_{t}}$, with $p_{i}$ prime and $e_{i}\geq 1$, of the orders of the cyclic groups q=#Gq=\#G.
Definition 3.2.
The order of an element in the cyclic group $\mathbb{F}_{p}^{\times}$ is defined by $\operatorname{ord}_{p}(v)=\min\{k:v^{k}\equiv 1\bmod p\}$. Primitive elements in this cyclic group have order p1=#Gp-1=\#G.
Lemma 3.2.
Let $G$ be a finite cyclic group of order p1=#Gp-1=\#G, and let $0\neq u\in G$ be an invertible element of the group. Then
[…]
The works in [9], and [46] attribute this formula to Vinogradov. The proof and other details on the characteristic function are given in [12, p. 863], [29, p. 258], [31, p. 18]. The characteristic function for multiple primitive roots is used in [7, p. 146] to study consecutive primitive roots. In [10] it is used to study the gap between primitive roots with respect to the Hamming metric. And in [46] it is used to prove the existence of primitive roots in certain small subsets $A\subset\mathbb{F}_{p}$. In [9] it is used to prove that some finite fields do not have primitive roots of the form $a\tau+b$, with $\tau$ primitive and $a,b\in\mathbb{F}_{p}$ constants. In addition, the Artin primitive root conjecture for polynomials over finite fields was proved in [38] using this formula.
3.3 Divisors Free Characteristic Function
It often difficult to derive any meaningful result using the usual divisors dependent characteristic function of primitive elements given in Lemma 3.2. This difficulty is due to the large number of terms that can be generated by the divisors, for example, $d\mid p-1$, involved in the calculations, see [12], [10] for typical applications and [30, p. 19] for a discussion.
A new divisors-free representation of the characteristic function of primitive element is developed here. This representation can overcomes some of the limitations of its counterpart in certain applications. The divisors representation of the characteristic function of primitive roots, Lemma 3.2, detects the order $\operatorname{ord}_{p}(u)$ of the element $u\in\mathbb{F}_{p}$ by means of the divisors of the totient p1p-1. In contrast, the divisors-free representation of the characteristic function, Lemma 3.3, detects the order $\operatorname{ord}_{p}(u)\geq 1$ of the element $u\in\mathbb{F}_{p}$ by means of the solutions of the equation $\tau^{n}-u=0$ in $\mathbb{F}_{p}$, where $u,\tau$ are constants, and $1\leq n<p-1,\gcd(n,p-1)=1,$ is a variable. Two versions are given: a multiplicative version, and an additive version.
Lemma 3.3.
Let $p\geq 2$ be a prime, and let $\tau$ be a primitive root mod $p$. For a nonzero element $u\in\mathbb{F}_{p}$, the followings hold:
(i) If $\psi\neq 1$ is a nonprincipal additive character of order $\operatorname{ord}\psi=p$, then
[…]
(ii) If $\chi\neq 1$ is a nonprincipal multiplicative character of order $\operatorname{ord}\chi=p-1$, then
[…]
where $\bar{u}$ is the inverse of $u\bmod p$.
Proof.
(i) As the index $n\geq 1$ ranges over the integers relatively prime to p1p-1, the element $\tau^{n}\in\mathbb{F}_{p}$ ranges over the primitive roots $\text{mod }p$. Ergo, the equation
$\tau^{n}-u=0$ (12)
has a solution if and only if the fixed element $u\in\mathbb{F}_{p}$ is a primitive root. Next, replace $\psi(z)=e^{i2\pi kz/p}$ to obtain
[…]
This follows from the geometric series identity $\sum_{0\leq k\leq N-1}w^{k}=(w^{N}-1)/(w-1)$ with $w\neq 1$, applied to the inner sum. For (ii), use the equation $\tau^{n}\bar{u}=1$, where $\bar{u}$ is the inverse of $u$, and apply the geometric series identity. ∎
4 Estimates Of Exponential Sums
This section provides simple estimates for the exponential sums of interest in this analysis. There are two objectives: To determine an upper bound, proved in Theorem 4.2, and to show that the finite Fourier transform
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\gcd(n,p-1)=1}e^{i2\pi\tau^{n}/p}+E(p),$ (14)
where $E(p)$ is an error term, this is proved in Lemma 4.1. These are indirectly implied by the equidistribution of the subsets
$\{\tau^{n}:\gcd(n,p-1)=1\}=\{b\tau^{n}:\gcd(n,p-1)=1\}\subset\mathbb{F}_{p},$ (15)
for any $0\neq b\in\mathbb{F}_{p}$. The proofs of these results are entirely based on established results and elementary techniques.
4.1 Incomplete And Complete Exponential Sums
Let $f:\mathbb{C}\longrightarrow\mathbb{C}$ be a function, and let $q\in\mathbb{N}$ be a large integer. The finite Fourier transform
$\hat{f}(t)=\frac{1}{q}\sum_{0\leq s\leq q-1}e^{i\pi st/q}$ (16)
and its inverse are used here to derive a summation kernel function, which is almost identical to the Dirichlet kernel.
Definition 4.1.
Let $p$ be a prime and let $\omega=e^{i2\pi/p}$ be a root of unity. The finite summation kernel is defined by the finite Fourier transform identity
$\mathcal{K}(f(n))=\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{0\leq s\leq p-1}\omega^{t(n-s)}f(s)=f(n).$ (17)
This simple identity is very effective in computing upper bounds of some exponential sums
$\sum_{n\leq x}f(n)=\sum_{n\leq x}\mathcal{K}(f(n)),$ (18)
where $x\leq p$. Two applications are illustrated here.
Theorem 4.1.
([44], [32]) Let $p\geq 2$ be a large prime, and let $\tau\in\mathbb{F}_{p}$ be an element of large multiplicative order $\operatorname{ord}_{p}(\tau)\mid p-1$. Then, for any $b\in[1,p-1]$, and $x\leq p-1$,
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}\ll p^{1/2}\log p.$ (19)
Proof.
Let $q=p+o(p)$ be a large prime, and let $f(n)=e^{i2\pi b\tau^{n}/p}$, where $\tau$ is a primitive root modulo $p$. Applying the finite summation kernel in Definition 4.1, yields
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}=\sum_{n\leq x}\frac{1}{q}\sum_{0\leq t\leq q-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}.$ (20)
The term $t=0$ contributes x/q-x/q, and rearranging it yield
$\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}$ $=$ $\frac{1}{q}\sum_{n\leq x,}\sum_{1\leq t\leq q-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{i2\pi b\tau^{s}/p}-\frac{x}{q}$
$=$ $\frac{1}{q}\sum_{1\leq t\leq q-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right)\left(\sum_{n\leq x}\omega^{tn}\right)-\frac{x}{q}.$
Taking absolute value, and applying Lemma 4.2, and Lemma 4.4, yield
$\left|\sum_{n\leq x}e^{i2\pi b\tau^{n}/p}\right|$ $\leq$ $\frac{1}{q}\sum_{1\leq t\leq q-1}\left|\sum_{0\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{n\leq x}\omega^{tn}\right|+\frac{x}{q}$
$\ll$ $\frac{1}{q}\sum_{1\leq t\leq q-1}\left(2q^{1/2}\log q\right)\cdot\left(\frac{2q}{\pi t}\right)+\frac{x}{q}$
$\ll$ $p^{1/2}\log^{2}p.$
The last summation in (4.1) uses the estimate
$\sum_{1\leq t\leq q-1}\frac{1}{t}\ll\log q\ll\log p$ (23)
since $q=p+o(p)$, and $x/q\leq 1$. ∎
This appears to be the best possible upper bound. The above proof generalizes the sum of resolvents method used in [32]. Here, it is reformulated as a finite Fourier transform method, which is applicable to a wide range of functions. A similar upper bound for composite moduli $p=m$ is also proved, [op. cit., equation (2.29)].
Theorem 4.2.
Let $p\geq 2$ be a large prime, and let $\tau$ be a primitive root modulo $p$. Then,
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi b\tau^{n}/p}\ll p^{1-\varepsilon}$ (24)
for any $b\in[1,p-1]$, and any arbitrary small number $\varepsilon\in(0,1/2)$.
Proof.
Let $p$ be a large prime, and let $f(n)=e^{i2\pi b\tau^{n}/p}$, where $\tau$ is a primitive root modulo $p$. Start with the representation
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}},$ (25)
see Definition 4.1. Use the inclusion exclusion principle to rewrite the exponential sum as
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=\sum_{1\leq n\leq p-1}\frac{1}{p}\sum_{0\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d).$ (26)
The term $t=0$ contributes $\varphi(p)/p$, and rearranging it yield
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
$=$ $\sum_{n\leq p-1}\frac{1}{p}\sum_{1\leq t\leq p-1,}\sum_{1\leq s\leq p-1}\omega^{t(n-s)}e^{\frac{i2\pi b\tau^{s}}{p}}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d)-\frac{\varphi(p)}{p}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p)}{p}.$
Taking absolute value, and applying Lemma 4.3, and Lemma 4.4, yield
$\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}\right|$
$\leq$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left|\sum_{1\leq s\leq p-1}\omega^{-ts}e^{i2\pi b\tau^{s}/p}\right|\cdot\left|\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right|+\frac{\varphi(p)}{p}$
$\ll$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(2qp{1/2}\log p\right)\cdot\left(\frac{4p^{1/2+\delta}}{\pi r(d,t)}\right)+\frac{\varphi(p)}{p}$
$\ll$ $p^{1/2+\delta}(\log p)^{2},$
where $r(d,t)\equiv dt\bmod p$ and $\delta>0$ is a small number. The last summation in (4.1) uses the estimate
$\sum_{1\leq t\leq p-1}\frac{1}{r(d,t)}=\sum_{1\leq z\leq p-1}\frac{1}{z}\ll\log p,$ (29)
since $z=r(d,t)\equiv dt\bmod p$ is a permutation in the finite field $\mathbb{F}_{p}$ and $\varphi(p)/p\leq 1$. This is restated in the simpler notation $p^{1/2+\delta}(\log p)^{2}\leq p^{1-\varepsilon}$ for any arbitrary small number $\varepsilon\in(0,1/2)$. ∎
The upper bound given in Theorem 4.2 seems to be optimum. A different proof, which has a weaker upper bound, appears in [15, Theorem 6], and related results are given in [6], [14], [20], and [21, Theorem 1].
4.2 Finite Fourier Transforms and Equivalent Exponential Sums
For any fixed $0\neq b\in\mathbb{F}_{p}$, the map $\tau^{n}\longrightarrow b\tau^{n}$ is one-to-one (permutation) in $\mathbb{F}_{p}$. Consequently, the subsets
$\{\tau^{n}:\gcd(n,p-1)=1\}\quad\text{ and }\quad\{b\tau^{n}:\gcd(n,p-1)=1\}\subset\mathbb{F}_{p}$ (30)
have the same cardinalities. As a direct consequence the exponential sums
$\sum_{\gcd(n,p-1)=1}e^{i2\pi b\tau^{n}/p}\quad\text{ and }\quad\sum_{\gcd(n,p-1)=1}e^{i2\pi\tau^{n}/p},$ (31)
have the same upper bound up to an error term. An asymptotic relation for the finite Fourier transform of exponential sums (31) is provided here.
Lemma 4.1.
Let $p\geq 2$ be a large prime. If $\tau$ be a primitive root modulo $p$ and $a\neq 0$ is not a primitive root, then
$\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}=-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}+O(p^{1/2+\delta}(\log p)^{2}),$ (32)
where $\delta>0$ is a small number and the implied constant is independent of $b\in[1,p-1]$.
Proof.
For $b\neq 1$, the exponential sum has the representation
$V(b)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p)}{p},$
confer equations (26) and (4.1) for more details. Similarly, for $b=1$,
$V(1)$ $=$ $\sum_{\gcd(n,p-1)=1}e^{\frac{i2\pi\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)-\frac{\varphi(p)}{p},$
respectively. Differencing (4.2) and (4.2) produces
$V(b)-V(1)$ $=$ $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)$
$\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right).$
Taking the finite Fourier transform of the difference D(a)=V(b)V(1)D(a)=V(b)-V(1) returns
$\widehat{D(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
$=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\times\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi b\tau^{s}}{p}}-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)$
$\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right)$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left(\sum_{1\leq s\leq p-1}\omega^{-ts}\sum_{1\leq b\leq p-1}e^{\frac{i2\pi b(\tau^{s}-a)}{p}}-\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)$
$\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right).$
By hypothesis $\tau^{s}-a\neq 0$ since $a\neq 0$ is not a primitive root. Thus, using the geometric sum identity $\sum_{1\leq u\leq p-1}e^{i2\pi u/p}=-1$ to simplify the last expression yields
$\widehat{D(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\left(\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}\right)$
$=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\left((-1)(-1)-(-1)\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)$
$\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right).$
Rearranging the last equation yield
$\widehat{V(a)}$ $=$ $\sum_{1\leq b\leq p-1}e^{-i2\pi\frac{ab}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi b\tau^{n}}{p}}$
$=$ $-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi\tau^{n}}{p}}+\frac{1}{p}\sum_{1\leq t\leq p-1}\left(1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\right)$
$\hskip 180.67499pt\times\left(\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\right).$
By Lemma 4.3, the relatively prime summation kernel is bounded by
$\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$ $=$ $\Bigg{|}\sum_{\gcd(n,p-1)=1}\omega^{tn}\Bigg{|}$ (39)
$\ll$ $\frac{p^{1+\delta}}{\pi r(d,t)},$
where $r(d,t)\equiv dt\bmod p$ and by Lemma 4.4, the difference including Gauss sum is bounded by
$\Bigg{|}1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg{|}=\Bigg{|}1-\sum_{1\leq s\leq p-1}\chi(s)\psi(s)\Bigg{|}$ $\leq$ $2p^{1/2}\log p,$ (40)
where $\chi(s)=e^{i\pi st/p}$, and $\psi(s)=e^{i2\pi\tau^{s}/p}$. Taking absolute value of the remainder term in (4.2) and replacing (39), and (40), return
$\widehat{R(a)}$ $=$ $\frac{1}{p}\sum_{1\leq t\leq p-1}\Bigg{|}1-\sum_{1\leq s\leq p-1}\omega^{-ts}e^{\frac{i2\pi\tau^{s}}{p}}\Bigg{|}\cdot\Bigg{|}\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1,\\ d\mid n\end{subarray}}\omega^{tn}\Bigg{|}$ (41)
$\ll$ $\frac{1}{p}(2p^{1/2}\log p)\cdot\sum_{1\leq t\leq p-1}\frac{p^{1+\delta}}{\pi r(d,t)}$
$\ll$ $\frac{1}{p}(2p^{1/2}\log p)\cdot\sum_{1\leq z\leq p-1}\frac{p^{1+\delta}}{\pi z}$
$\ll$ $p^{1/2+\delta}(\log p)^{2},$
since for each $d\mid p-1$, the map $t\longrightarrow z=r(d,t)\equiv dt\bmod p$ permutes the elements in $\mathbb{F}_{p}$. where the implied constant depends on the number of divisors of p1p-1. ∎
The same proof works for many other subsets of elements $\mathcal{A}\subset\mathbb{F}_{p}$. For example,
$\sum_{n\in\mathcal{A}}e^{i2\pi b\tau^{n}/p}=\sum_{n\in\mathcal{A}}e^{i2\pi\tau^{n}/p}+O(p^{1/2+\delta}\log^{c}p),$ (42)
for some constant $c>0$.
4.3 Finite Summation Kernels And Gaussian Sums
Lemma 4.2.
Let $p\geq 2$ be large prime and let $\omega=e^{i2\pi/p}$ be a $p$th root of unity, and let $t\in[1,p-1]$. Then,
  1. (i)
    $\sum_{n\leq p-1}\omega^{tn}=\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}},$
  2. (ii)
    $\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}\leq\frac{2p}{\pi t}.$
Proof.
(i) Use the geometric series to compute this simple exponential sum as
$\sum_{n\leq p-1}\omega^{tn}$ $=$ $\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}.$
(ii) Observe that the parameters are $p$ prime, $\omega=e^{i2\pi/p}$, the integers $t\in[1,p-1]$, and $d\leq p-1<q-1$. This data implies that $\pi t/q\neq k\pi$ with $k\in\mathbb{Z}$, so the sine function $\sin(\pi t/p)\neq 0$ is well defined. Using standard manipulations, and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$, the last expression becomes
$\bigg{|}\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}\bigg{|}\leq\bigg{|}\frac{2}{\sin(\pi t/p)}\bigg{|}\leq\frac{2p}{\pi t}.$ (43)
Lemma 4.3.
Let $p\geq 2$ be large prime, and let $\omega=e^{i2\pi/p}$ be a $p$th root of unity. Then,
  1. (i)
    $\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}=\sum_{d\mid p-1}\mu(d)\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}},$
  2. (ii)
    $\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\right|\ll\frac{p^{1+\delta}}{\pi r(d,t)},$
    where $r(d,t)\equiv dt\bmod p$ is a permutation of $t\in\mathbb{F}_{p}$ and $\delta>0$ is a small number, for any $t\in[1,p-1]$.
Proof.
(i) Use the inclusion exclusion principle to rewrite the exponential sum as
$\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}$ $=$ $\sum_{n\leq p-1}\omega^{tn}\sum_{\begin{subarray}{c}d\mid p-1\\ d\mid n\end{subarray}}\mu(d)$
$=$ $\sum_{d\mid p-1}\mu(d)\sum_{\begin{subarray}{c}n\leq p-1\\ d\mid n\end{subarray}}\omega^{tn}$
$=$ $\sum_{d\mid p-1}\mu(d)\sum_{m\leq(p-1)/d}\omega^{dtm}$
$=$ $\sum_{d\mid p-1}\mu(d)\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}}.$
(ii) Observe that the parameters are $p$ prime, $\omega=e^{i2\pi/p}$, the integers $t\in[1,p-1]$ and $d\leq p-1<p$. In addition, the map $t\longrightarrow dt$ is a permutation in the finite field $\mathbb{F}_{p}$ for any fixed $d\neq 0$. This data implies that $\pi dt/p\neq k\pi$ with $k\in\mathbb{Z}$, so the sine function $\sin(\pi dt/p)\neq 0$ is well defined. Using standard manipulations, and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$, the last expression becomes
$\left|\frac{\omega^{dt}-\omega^{dtp}}{1-\omega^{dt}}\right|\leq\left|\frac{2}{\sin(\pi dt/p)}\right|\leq\frac{2p}{\pi r(d,t)},$ (45)
where $r(d,t)\equiv dt\bmod p$. Finally, the upper bound is
$\left|\sum_{d\mid p-1}\mu(d)\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}}\right|$ $\leq$ $\sum_{d\mid p-1}\left|\frac{2}{\sin(\pi dt/p)}\right|$ (46)
$\leq$ $\sum_{d\mid p-1}\frac{2p}{\pi r(d,t)}$
$\ll$ $\frac{p^{1+\delta}}{\pi r(d,t)},$
where $\sum_{d\mid p-1}1=d(p-1)\ll p^{\delta}$ is the number of divisor in p1p-1 and $\delta>0$ is a small number. ∎
Lemma 4.4.
(Gauss sums) Let $p\geq 2$ and $q$ be large primes. Let $\chi(t)=e^{i2\pi t/q}$ and $\psi(t)=e^{i2\pi\tau^{t}/p}$ be a pair of characters. Then, the Gaussian sum has the upper bound
$\left|\sum_{1\leq t\leq q-1}\chi(t)\psi(t)\right|\leq 2q^{1/2}\log q.$ (48)
5 Evaluation Of The Main Term
Finite sums and products over the primes numbers occur on various problems concerned with primitive roots. These sums and products often involve the normalized totient function $\varphi(n)/n=\prod_{p|n}(1-1/p)$ and the corresponding estimates, and the asymptotic formulas.
Lemma 5.1.
([43, Lemma 1]) Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (49)
where $\operatorname{li}(x)$ is the logarithm integral, $a_{1}=0.373955\ldots$, and $B>1$ is an arbitrary constant, as $x\rightarrow\infty$.
A more general version of this Lemma is proved in [45], and related discussions are given in [30, p. 16]. The generalization to number fields appears in [24]. These results are ubiquitous in various results in Number Theory. Here, the logarithm integral is defined by
$\operatorname{li}(x)=\int_{2}^{x}\frac{1}{\log t}dt=\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right).$ (50)
In this application, the constant
$a_{1}=\prod_{p\geq 2}\left(1-\frac{1}{p(p-1)}\right)$ (51)
coincides to the average density of primitive roots modulo $p$. The average density of primitive roots was proved in [18] and [43], (the heuristic is due to Artin, and the average density is known as Artin constant).
Lemma 5.2.
Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1=a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right).$ (52)
Proof.
A routine rearrangement gives
$\sum_{p\leq x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1$ $=$ $\sum_{p\leq x}\frac{\varphi(p-1)}{p}$
$=$ $\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}.$
To proceed, apply Lemma 5.1 to reach
$\sum_{p\leq x}\frac{\varphi(p-1)}{p-1}-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ $=$ $a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right)-\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}$ (54)
$=$ $a_{1}\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$
where the second finite sum
$\sum_{p\leq x}\frac{\varphi(p-1)}{p(p-1)}=O\left(\log\log x\right)$ (55)
is absorbed into the error term, which has $B>1$ as an arbitrary constant. ∎
6 Estimate For The Error Term
The upper bounds for exponential sums over subsets of elements in finite rings $\mathbb{Z}/m\mathbb{Z}$ or finite fields $\mathbb{F}_{p}$ studied in Lemma 4.1 and Theorem 4.2 are used to estimate the error term $E(x)$ in the proof of Theorem 1.1.
Lemma 6.1.
Let $p\geq 2$ be a large prime, let $\psi\neq 1$ be an additive character, and let $\tau$ be a primitive root mod $p$. Assume the element $u\neq 0$ is not a primitive root. Then,
$\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\sum_{0<k\leq p-1}\psi\left((\tau^{n}-u)k\right)\ll\frac{x^{1-\varepsilon}}{\log x}$ (56)
for all sufficiently large numbers $x\geq 1$ and an arbitrarily small number $\varepsilon<1/16$.
Proof.
Let $\psi(z)=e^{i2\pi z/p}$. By hypothesis $u\neq\tau^{n}$ for any $n\geq 1$ such that $\gcd(n,p-1)=1$. Rearrange the triple finite sum in the form
$E(x)$ $=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{0<k\leq p-1,}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\psi((\tau^{n}-u)k)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\left(\sum_{0<k\leq p-1}e^{-i2\pi\frac{uk}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{k\tau^{n}}{p}}\right)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\left(-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{\tau^{n}}{p}}+O(p^{1/2+\delta}\log^{2}p)\right).$
The last line follows from Lemma 4.1. By Theorem 4.2, the inner sum has the upper bound
$\bigg{|}\widehat{V(u)}\bigg{|}$ $=$ $\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\tau^{n}/p}+O\left(p^{1/2+\delta}(\log p)^{2}\right)\right|$
$\ll$ $\left|\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\tau^{n}/p}\right|+p^{1/2+\delta}(\log p)^{2}$
$\ll$ $p^{1-\varepsilon},$
where $\varepsilon<1/2$ is an arbitrarily small number. Replacing the estimate (6) back into (6) and completing the finite sum return
$|E(x)|$ $\ll$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\cdot p^{1-\varepsilon}$
$\ll$ $\frac{1}{x^{\varepsilon}}\sum_{x\leq p\leq 2x}1$
$\ll$ $\frac{x^{1-\varepsilon}}{\log x},$
where the number of primes in the short interval $[x,2x]$ is $\pi(2x)-\pi(x)\leq 2x/\log x$. ∎
7 Densities Of Primes And Primitive Roots
The representations of the characteristic function of primitive roots, Lemma 3.2, and Lemma 3.3, easily detect certain local and global properties of the elements $u\in\mathbb{F}_{p}$ in a finite field. Exempli gratia, it vanishes
[…]
Ergo, the constraints $u\neq\pm 1,v^{2}$ with $v\in\mathbb{Z}$, are necessary global constraints to be a primitive element in $\mathbb{F}_{p}$, $p\geq 2$ an arbitrary prime. But the constraints $u\neq v^{q},q\in\mathcal{Q}$, where $\mathcal{Q}$ is a finite subset of primes, are not necessary global constraints since there are infinitely many primes for which $q\nmid p-1,q\in\mathcal{Q}$. The requirement of being dth power nonresidues $\bmod p$ for all $d\mid p-1$, which is equivalent to the definition of primitive root, are necessary local properties.
Proof.
(Theorem 1.1): Suppose that $u\neq\pm 1,v^{2}$ is not a primitive root for all primes $p\geq x_{0}$, with $x_{0}\geq 1$ constant. Let $x>x_{0}$ be a large number, and consider the sum of the characteristic function over the short interval $[x,2x]$, that is,
$0=\sum_{x\leq p\leq 2x}\Psi(u).$ (61)
Replacing the characteristic function, Lemma 3.3, and expanding the nonexistence equation (61) yield
$0$ $=$ $\sum_{x\leq p\leq 2x}\Psi(u)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\gcd(n,p-1)=1,}\sum_{0\leq k\leq p-1}\psi\left((\tau^{n}-u)k\right)$
$=$ $\delta(u)\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\gcd(n,p-1)=1}1+\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\gcd(n,p-1)=1,}\sum_{0<k\leq p-1}\psi\left((\tau^{n}-u)k\right)$
$=$ $\delta(u)M(x)+E(x),$
where $\delta(u)\geq 0$ is a constant depending on the fixed integer $u\neq 0$.
The main term $M(x)$ is determined by a finite sum over the trivial additive character $\psi=1$, and the error term $E(x)$ is determined by a finite sum over the nontrivial additive characters $\psi(t)=e^{i2\pi t/p}\neq 1$.
Applying Lemma 5.2 to the main term, and Lemma 6.1 to the error term yield
$\sum_{x\leq p\leq 2x}\Psi(u)$ $=$ $\delta(u)M(x)+E(x)$
$=$ $\delta(u)\left(\text{li}(2x)-\text{li}(x)\right)+O\left(\frac{x}{\log^{B}x}\right)+O\left(\frac{x^{1-\varepsilon}}{\log x}\right)$
$=$ $\delta(u)\frac{x}{\log x}+O\left(\frac{x}{\log^{B}x}\right)$
$>$ $0,$
where $B>1$. However, $\delta(u)>0$ contradicts the hypothesis (61) for all sufficiently large numbers $x\geq x_{0}$. Ergo, the short interval $[x,2x]$ contains primes with the fixed primitive root $u$. ∎
A formula for computing the density constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is derived in [23, p. 220]. This formula specifies the density as follows. Let $u=(st^{2})^{k}\neq\pm 1,v^{2}$ with $s$ squarefree, and $k\geq 1$, and let
$a_{k}(u)=\prod_{p\mid k}\frac{1}{p-1}\prod_{p\nmid k}\left(1-\frac{1}{p(p-1)}\right).$ (64)
  1. 1.
    If $s\not\equiv 1\bmod 4$, then,
    $\delta(u)=a_{k}(u).$ (65)
  2. 2.
    If $s\equiv 1\bmod 4$, then,
    $\delta(u)=\left(1-\mu(s)\prod_{\begin{subarray}{c}p\mid s\\ p\mid k\end{subarray}}\frac{1}{p-2}\prod_{\begin{subarray}{c}p\mid s\\ p\nmid k\end{subarray}}\frac{1}{p^{2}-p-1}\right)a_{k}(u).$ (66)
In the case of a decimal base $u=10$, the density (51), which is computed in [47], is
$\delta(10)=\prod_{p\geq 2}\left(1-\frac{1}{p(p-1)}\right)=0.37739558136192022880547280\ldots.$ (67)
The argument can be repeated for any other squarefree base $u=2,3,\ldots$.
Information on the determination of the constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is provided in the next section and in [23, p. 220].
8 Application To Repeated Decimals
Let $p\geq 2$ be a prime. The period of the repeating decimal number
$1/p=0.\overline{x_{d-1}\ldots x_{1}x_{0}},$ (68)
with $x_{i}\in\{0,1,2,\ldots,9\}$, was investigated by Gauss and earlier authors centuries ago, see [5] for a historical account, and [33] for recent developments. As discussed in Articles 14-18 in [17], the period, denoted by $\operatorname{ord}_{p}(10)=d\geq 1$, is a divisor of p1p-1. The problem of computing the densities for the subsets of primes for which the repeating decimals have very large periods such as d=(p1)/2d=(p-1)/2, and d=p1d=p-1, is a recent problem. This note considers the following result.
Theorem 8.1.
There are infinitely many primes $p\geq 7$ with maximal repeating decimal $1/p=0.\overline{x_{p-2}x_{p-3}\cdots x_{1}x_{0}}$, where $0\leq x_{i}\leq 9$. Moreover, the counting function for these primes satisfies the lower bound
$\pi_{10}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(10)=p-1\right\}=\delta(10)\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right),$ (69)
where $\delta(10)=0.3773955\ldots$ is a constant, see (67), for all large numbers $x\geq 1$.
The proof of Theorem 8.1 is a corollary of the more general result in Theorem 1.1 presented in Section 7. This analysis generalizes to repeating $\ell$-adic expansions $1/p=0.\overline{x_{d-1}x_{d-2}\cdots x_{1}x_{0}}$, where $0\leq x_{i}\leq\ell-1$, in any numbers system with nonsquare integer base $\ell\geq 2$.
9 Proof of Theorem 8.1
The repeating decimal fractions have the squarefree base $u=10$. In particular, the repeated fraction representation
$\frac{1}{p}=\frac{m}{10^{d}}+\frac{m}{10^{2d}}+\cdots=m\sum_{n\geq 1}\frac{1}{10^{dn}}=\frac{m}{10^{d}-1}$ (70)
has the maximal period d=p1d=p-1 if and only if 10 has order $\operatorname{ord}_{p}(10)=p-1$ modulo $p$. This follows from the Fermat little theorem and $10^{d}-1=mp$. The subset of Abel-Wieferich primes, which satisfy the congruence
$10^{p-1}-1\equiv 0\bmod p^{2},$ (71)
are also important in the calculation of the period modulo $p^{2}$, confer [39, p. 333], [11] for other details.
The result for repeating decimal of maximal period is a simple corollary of the previous result.
Proof.
(Theorem 8.1) This follows from Theorem 1.1 replacing the base $u=10$. That is,
$\pi_{10}(x)=\sum_{p\leq x}\Psi(10)=\delta(10)\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (72)
where $B>1$ is a constant. ∎